Give an example for each of the following.

i) A set in R with a unique limit point.

ii) A set in R whose all points except the one are its limit points.

iii) A set having no limit point.

iv) A set S with S°= S̅

v) A bijection from Nodd to Z


1
Expert's answer
2021-03-30T06:47:29-0400

i) "\{\frac{1}{n}:n\in \mathbb{N}\}\cup\{0\}" has 0 as only limit point.

"ii)[0,1]\cup \{2\}" has all points except 2 as its limit point.

iii) "\{1,2\}" has no limit point.

iv) S="(-1,1)" the open interval as a subset of "(-1,1)\cup (5,\infty)" in subspace topology.

v)"f:\mathbb{N}_{odd}\longrightarrow \mathbb{Z}" is given by

"8n+1\mapsto 2n" ,

"8n+3\mapsto -2n" ,

"8n+5\mapsto 2n+1,"

"8n+7\mapsto -2n+1." This clearly a bijection.